# Ranks of matrix factorizations and sheaf cohomology

### v1

@@ 1. Introduction
We will use the following notation throughout the paper. Let $k$ be a field, $S\coloneqq k[x_{0},\dots,x_{n}]$ a standard graded polynomial ring, $f\in S$ a homogeneous and irreducible polynomial of degree $d$, $R\coloneqq S/(f)$, ${\mathfrak{m}}$ the homogeneous maximal ideal of $R$, $X$ the projective hypersurface $\operatorname{Proj}(R)$, and $i$ the closed embedding $X\hookrightarrow\mathbb{P}^{n}_{k}$. We assume $n\geq 1$ throughout (otherwise $X$ is empty). We write $a\coloneqq n+1-d$ and $e\coloneqq\lfloor\frac{n}{2}\rfloor$.
@@ 2.1. Background on graded matrix factorizations
[…]
@@ 2.2. Orlov’s Theorem
[…]
@@ 2.3. Graded Auslander duality
Let $A=\bigoplus_{i\geq 0}A_{i}$ be a commutative, graded Gorenstein ring of dimension $t$ such that $A_{0}=k$, and assume $A$ has an isolated singularity. There is an isomorphism $\mathbf{R}{\operatorname{\underline{Hom}}}_{A}(k,A)\cong k(b)[-t]$ in $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(A)$ for some $b,t\in\mathbb{Z}$. For example, our graded hypersurface ring $R$ is Gorenstein, and in this case $b$ is equal to $a=n+1-d$, and $t$ is equal to $n$. The following result is a version of Auslander duality [Aus78] for the graded singularity category $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$ (see also [Mur13, (3.7)]). Given a graded $k$-vector space $V$, let $V^{*}\coloneqq\operatorname{\underline{Hom}}_{k}(V,k)$ denote its $k$-dual.
@@
We have $F[2]=F(d)$ for all $F\in\operatorname{hmf}_{\operatorname{gr}}(f)$, and the equivalence in (2.1) commutes with grading twists and cohomological shifts; it follows that $k(\ell)[2j]\cong k(\ell+jd)$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ for all $j\in\mathbb{Z}$. Since the functor $\Phi_{0}$ commutes with cohomological shifts, we may assume $q=0$, i.e. $-d<\ell\leq 0$. We therefore must show $\Phi_{0}(k(\ell))\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}}(\ell-a))[n+\ell-a-1]$.
@@
on the variables $x_{0},\dots,x_{n}\in S$. The underlying module of $F$ is thus $K\otimes_{S}D\otimes_{S}R$, where $D$ is a divided power algebra on a single variable $t$ of cohomological degree $-2$ and internal degree $d$. We therefore have:
@@
where $R(\ell-a)$ is in cohomological degree $-n$. Observe that
@@
$\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\succeq 0},\$
where $R(\ell-a)$ is in cohomological degree $-n+1$. Notice that $\ell-a+d>0$ (this is where we use the assumption $a\leq 0$). Thus, if $\ell>a$, then $\Phi_{0}(k(\ell))=0$; otherwise, $\Phi_{0}(k(\ell))$ is isomorphic to
@@
a brutally truncated Koszul complex on $x_{0},\dots,x_{n}$, where $\mathcal{O}_{X}(\ell-a)$ is in cohomological degree $-n+1$. This complex is a locally free resolution of the sheaf $i^{*}\operatorname{ker}\left(\mathcal{O}_{\mathbb{P}^{n}_{k}}(1)^{\binom{n+1}{a-\ell+1}}\to\mathcal{O}_{\mathbb{P}^{n}_{k}}(2)^{\binom{n+1}{a-\ell+2}}\right)\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}})(\ell-a)$ in cohomological degree $a-\ell-n+1$, where the isomorphism follows from [Eis95, Theorem 17.16]. ∎
@@
    Fix $i\in\{0,1\}$ and $j\in\mathbb{Z}$, and write $a-j=qd-r$ for $0\leq r<d$. We have:
@@
Thus, the graded Betti number of $M$ in cohomological degree 0 (resp. $-1$) and internal degree $j$ is $b^{1}_{j}$ (resp. $b^{0}_{j})$. We now compute:
$b^{i}_{j}$ $=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(\Psi_{a}(\mathcal{C}),k(-j)[i-1]\right)$
$=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(k(a-j),\Psi_{a}(\mathcal{C})[n-i]\right)$
$=\dim_{k}\operatorname{Ext}^{n-i}_{X}\left(\Phi_{0}(k(a-j)),\mathcal{C}\right)$
$=\dim_{k}\operatorname{Ext}_{X}^{r+a+1-i-2q}\left(i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a),\mathcal{C}\right)$
$=h^{r+a+1-i-2q}\left(X,\mathcal{C}\otimes i^{*}(\Omega^{r+a}_{\mathbb{P}^{n}_{k}})(r+a)\right)$
@@
where the first equality follows from [Pav21, Lemma 3.1] and the identification of the shift $M[1]$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ with the first syzygy of $M$, the second from Proposition 2.7, the third from Proposition 2.6, the fourth from Theorem 3.1, the fifth from a straightforward calculation, and the sixth from the projection formula. This proves (1). For (2), we have $\operatorname{rank}(F)=\operatorname{rank}(F^{0})+\operatorname{rank}(F^{1})=\sum_{j\in\mathbb{Z}}(b^{0}_{j}+b^{1}_{j})~{}=\rho(\mathcal{C}).$
@@ 4. Examples
Let $\mathcal{F}$ be the structure sheaf of a point on $X$. We evidently have
$\rho(\mathcal{F})=\sum_{r=0}^{n}\operatorname{rank}(\Omega^{r}_{\mathbb{P}^{n}_{k}})=2^{n+1}-1>2^{e+1}.$
@@
(4.5) $0\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-d)\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to\mathcal{O}_{X}\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to 0$
@@
Assume $k$ contains a square root $i$ of $-1$, and suppose $f=x_{0}^{n}+\cdots+x_{n}^{n}$, where $n$ is even. In this case, $e=\frac{n}{2}$. We have graded matrix factorizations

### v2

@@ 1. Introduction
We will use the following notation throughout the paper. Let $k$ be a field, $S\coloneqq k[x_{0},\dots,x_{n}]$ a standard graded polynomial ring, $f\in S$ a homogeneous and irreducible polynomial of degree $d$, $R\coloneqq S/(f)$, $\mathfrak{m}$} the homogeneous maximal ideal of $R$, $X$ the projective hypersurface $\operatorname{Proj}(R)$, and $i$ the closed embedding $X\hookrightarrow\mathbb{P}^{n}_{k}$. We assume $n\geq 1$ throughout. We write $a\coloneqq n+1-d$ and $e\coloneqq\lfloor\frac{n}{2}\rfloor$.
@@ 2.1. Background on graded matrix factorizations
${F^{1}(-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s^{1}}$$\scriptstyle{\alpha^{1}}$${F^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s^{0}}$$\scriptstyle{\alpha^{0}}$${F^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha^{1}}$${G^{1}(-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t^{1}}$${G^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t^{0}}$${G^{1}}$
@@ 2.2. Orlov’s Theorem
Suppose $M$ is an MCM module generated in degree 0; we will discuss explicit examples of such modules in Example 4.7 below. We now show $\Phi_{0}(M)=\widetilde{M}$. Let $F$ be the minimal free resolution of $M$, which has the form $F=\left[P\xleftarrow{\alpha}Q\xleftarrow{\beta}P(-d)\xleftarrow{\alpha}Q(-d)\xleftarrow{\beta}\cdots\right]$ for some graded free $R$-modules $P$ and $Q$ and maps $\alpha$ and $\beta$, where $P$ is a direct sum of copies of $R$. Given a graded $R$-module NN, write $N^{\vee}\coloneqq\operatorname{\underline{Hom}}_{R}(N,R)$. We have $F_{\succeq 0}=F$, and $\operatorname{\underline{Hom}}_{R}(F,R)=\left[P^{\vee}\xrightarrow{\alpha^{\vee}}Q^{\vee}\xrightarrow{\beta^{\vee}}P^{\vee}(d)\xrightarrow{\alpha^{\vee}}Q^{\vee}(d)\xrightarrow{\beta^{\vee}}\cdots\right]$. The minimal free resolution of $\operatorname{\underline{Hom}}_{R}(F,R)$ is the complex $F^{\prime}=\left[Q^{\vee}(-d)\xleftarrow{\alpha^{\vee}}P^{\vee}(-d)\xleftarrow{\beta^{\vee}}Q^{\vee}(-2d)\xleftarrow{\alpha^{\vee}}P^{\vee}(-2d)\xleftarrow{\beta^{\vee}}\cdots\right]$. Dualizing $F^{\prime}$ gives the complex $\left[Q(d)\xrightarrow{\alpha}P(d)\xrightarrow{\beta}Q(2d)\xrightarrow{\alpha}P(2d)\xrightarrow{\beta}\cdots\right]$, whose terms are direct sums of copies of $R(j)$ for various $j>0$. We conclude that $\varphi_{0}(M)=\operatorname{\underline{Hom}}_{R}(F^{\prime},R)$, which is quasi-isomorphic to $M$. Thus, $\Phi_{0}(M)=\widetilde{M}$.
@@ 2.3. Graded Auslander duality
Let $A=\bigoplus_{i\geq 0}A_{i}$ be a commutative, graded Gorenstein ring of dimension $t$ such that $A_{0}=k$, and assume $A$ has an isolated singularity. There is an isomorphism $\mathbf{R}{\operatorname{\underline{Hom}}}_{A}(k,A)\cong k(b)[-t]$ in $\operatorname{D}^{\operatorname{b}}_{\operatorname{gr}}(A)$ for some $b,t\in\mathbb{Z}$. For example, our graded hypersurface ring $R$ is Gorenstein, and in this case $b$ is equal to a=n+1da=n+1-d, and $t$ is equal to $n$. The following result is a version of Auslander duality [Aus78] for the graded singularity category $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(A)$ (see also [Mur13, (3.7)]). Given a graded $k$-vector space $V$, let $V^{*}\coloneqq\operatorname{\underline{Hom}}_{k}(V,k)$ denote its $k$-dual.
@@
We have $F[2]=F(d)$ for all $F\in\operatorname{hmf}_{\operatorname{gr}}(f)$, and the equivalence in (2.1) commutes with grading twists and cohomological shifts; it follows that $k(\ell)[2j]\cong k(\ell+jd)$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ for all $j\in\mathbb{Z}$. Since the functor $\Phi_{0}$ commutes with cohomological shifts, we may assume $q=0$, i.e. $d<\ell\leq 0$. We therefore must show $\Phi_{0}(k(\ell))\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}}(\ell-a))[n+\ell-a-1]$.
@@
on the variables $x_{0},\dots,x_{n}\in S$. The underlying module of $F$ is thus $K\otimes_{S}D\otimes_{S}R$, where $D$ is a divided power algebra on a single variable $t$ of cohomological degree 2-2 and internal degree $d$. We therefore have:
@@
where $R(\ell-a)$ is in cohomological degree n-n. Observe that
@@
$\left[R(\ell-a)\to R(\ell-a+1)^{n+1}\to R(\ell-a+2)^{\binom{n+1}{2}}\oplus R(\ell-a+d)\to\cdots\right]_{\succeq 0},\\$
where $R(\ell-a)$ is in cohomological degree n+1-n+1. Notice that $\ell-a+d>0$ (this is where we use the assumption $a\leq 0$). Thus, if $\ell>a$, then $\Phi_{0}(k(\ell))=0$; otherwise, $\Phi_{0}(k(\ell))$ is isomorphic to
@@
a brutally truncated Koszul complex on $x_{0},\dots,x_{n}$, where $\mathcal{O}_{X}(\ell-a)$ is in cohomological degree n+1-n+1. This complex is a locally free resolution of the sheaf $i^{*}\operatorname{ker}\left(\mathcal{O}_{\mathbb{P}^{n}_{k}}(1)^{\binom{n+1}{a-\ell+1}}\to\mathcal{O}_{\mathbb{P}^{n}_{k}}(2)^{\binom{n+1}{a-\ell+2}}\right)\cong i^{*}(\bigwedge^{a-\ell}\mathcal{T}_{\mathbb{P}^{n}_{k}})(\ell-a)$ in cohomological degree $a-\ell-n+1$, where the isomorphism follows from [Eis95, Theorem 17.16]. ∎
@@
    Fix $i\in\{0,1\}$ and $j\in\mathbb{Z}$, and write aj=qdra-j=qd-r for $0\leq r<d$. We have:
@@
Thus, the graded Betti number of $M$ in cohomological degree 0 (resp. 1-1) and internal degree $j$ is $b^{1}_{j}$ (resp. $b^{0}_{j})$. By [Pav21, Lemma 3.1] and the identification of the shift $M[1]$ in $\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)$ with the first syzygy of $M$, we have $b^{i}_{j}=\dim_{k}\operatorname{Hom}_{\operatorname{D}^{\operatorname{\operatorname{sing}}}_{\operatorname{gr}}(R)}\left(\Psi_{a}(\mathcal{C}),k(-j)[i-1]\right)$. We now compute:
[…]
Finally, we have:
$\operatorname{Ext}_{X}^{r+a+1-i-2q}\left(i^{*}(\bigwedge^{r+a}\mathcal{T}_{\mathbb{P}^{n}_{k}})(-r-a),\mathcal{C}\right)$ $=H^{r+a+1-i-2q}\left(X,\mathcal{C}\otimes i^{*}(\Omega^{r+a}_{\mathbb{P}^{n}_{k}})(r+a)\right)$
@@
where the second equality follows from the projection formula. This proves (1). For (2), we note that $\operatorname{rank}(F)=\operatorname{rank}(F^{0})+\operatorname{rank}(F^{1})=\sum_{j\in\mathbb{Z}}(b^{0}_{j}+b^{1}_{j})~=\rho(\mathcal{C}).$
@@ 4. Examples
Let $\mathcal{F}$ be the structure sheaf of a point on $X$, and assume $n\geq 1$. We evidently have
$\rho(\mathcal{F})=\sum_{r=0}^{n}\operatorname{rank}(\Omega^{r}_{\mathbb{P}^{n}_{k}})=2^{n}\geq 2^{e+1}.$
@@
(4.5) $0\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-d)\to\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to\mathcal{O}_{X}\otimes\Omega^{r}_{\mathbb{P}^{n}_{k}}(r)\to 0.$
@@
As a concrete example, take $n=2$, j=1j=-1, and $d=1$. Formula 4.4 implies that
$h^{i}(\mathbb{P}^{n}_{k},\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-2))=\begin{cases}1,&r=i=2,\\ 0,&\text{else};\end{cases}\quad\text{and}\quad h^{i}(\mathbb{P}^{n}_{k},\Omega^{r}_{\mathbb{P}^{n}_{k}}(r-1))=\begin{cases}1,&r=i=1,\\ 0,&\text{else}.\end{cases}$
It thus follows from the short exact sequence (4.5) (twisted by 1-1) that
$\rho(\mathcal{O}_{X}(-1))=h^{1}(\mathbb{P}^{n}_{k},\mathcal{O}_{X}\otimes\Omega^{1}_{\mathbb{P}^{n}_{k}})+h^{1}(\mathbb{P}^{n}_{k},\mathcal{O}_{X}\otimes\Omega^{2}_{\mathbb{P}^{n}_{k}}(1))=h^{1}(\mathbb{P}^{n}_{k},\Omega^{1}_{\mathbb{P}^{n}_{k}})+h^{2}(\mathbb{P}^{n}_{k},\Omega^{2}_{\mathbb{P}^{n}_{k}})=2.$
A nearly identical calculation shows that, when $n=d=2$, we have $\rho(\mathcal{O}_{X})=2$.
@@
Assume $k$ contains a square root $i$ of 1-1, and suppose $f=x_{0}^{n}+\cdots+x_{n}^{n}$, where $n$ is even. In this case, $e=\frac{n}{2}$. We have graded matrix factorizations